OpenStudy (anonymous):

check all of the order pairs that satisfy the equation below. y= 2/3x a. (21,3/2) B. (7,14) c. (21,14) D. (28,42) E. (42,28) F (14,21)

OpenStudy (anonymous):

helppp -.-

OpenStudy (mathstudent55):

Take each ordered pair and plug it into the equation. If it makes the equation true, then it satisfies the equation. I'll use the first point as an example. Equation y = (2/3)x a. Test (21, 3/2) in the equation Is 3/2 = (2/3)*21? 3/2 = 14 is false, so (21, 3/2) does not satisfy the equation.

OpenStudy (anonymous):

ok i get it but at the same time i dont so A. i know that one does not satisfy the equation but ok like i said still kinda lost

OpenStudy (anonymous):

nevermind i got the answer i get now thanks for your help though (:

OpenStudy (mathstudent55):

Let's do the third together as another example. Equation: y = (2/3)x c. Test (21, 14) We replace x in the equation with 21, and we replace y in the equation with 14. Is 14 = (2/3) * 21? 14 = 14 True So (21, 14) satisfies the equation.

OpenStudy (mathstudent55):

wlcm